OpenStudy (anonymous):

Help Proof Reading my Algebra 2 essay Question? To calculate the volume of a chemical produced in a day a chemical manufacturing company uses the following formula below: V(x) = [C1(x) + C2(x)](H(x)) where x represents the number of units produced. This means two chemicals are added together to make a new chemical and the resulting chemical is multiplied by the expression for the holding container with respect to the number of units produced. The equations for the two chemicals added together with respect to the number of unit produced are given below: C1(x) = x/x+1,C2(x) = 2/x-3

OpenStudy (anonymous):

The equation for the holding container with respect to the number of unit produced is given below: H(x) = x^3 - 9x/x a. What rational expression C(x) do you get when you combine the two chemicals? b. What is the simplified equation of V(x)? c. What would the volume be if 50, 100, or 1000 units are produced in a day? d. The company needs a volume of 3000 c^3 How many units would need to be produced in a day

OpenStudy (anonymous):

I got To add fractions, give them a common denominator. 2/3 + 5/7 = 14/21 + 15/21 = 29/21 x/(x+1) + 2/(x-3) = x(x-3)/((x+1)(x-3)) + 2(x+1)/((x+1)(x-3)) = (x^2-x+2)/((x+1)(x-3)) b) The factor (x-3) cancels. c) and d) are just finding V from a value for x and x from a value for V. can you help me make this look better for an answer?